OpenStudy (anonymous):

A line has a slope of 9/4 and a y-intercept of ���������14. Which equations represent the line? Choose all answers that are correct.

OpenStudy (anonymous):

@animal_lover36

OpenStudy (anonymous):

@CausticSyndicalist

OpenStudy (anonymous):

@Destinycholie

OpenStudy (anonymous):

@dragonmaster917

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@monsteygoo

OpenStudy (anonymous):

@One098

OpenStudy (anonymous):

@Rainbow_dashie

OpenStudy (shinalcantara):

slope-intercept form: y= mx + b y = (9/4)x -14 multiplying the whole equation with 4 you'll have: 4y = 9x - 14(4) 4y - 9x + 56 = 0

OpenStudy (anonymous):

@shortycme

OpenStudy (anonymous):

shina is correct

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

sorry i was doing something and i also lost connection :P

OpenStudy (anonymous):

but wait

OpenStudy (rainbow_dashie):

Calm yourself

OpenStudy (rainbow_dashie):

let me try too help

OpenStudy (anonymous):

@shinalcantara

OpenStudy (rainbow_dashie):

:)

OpenStudy (anonymous):

and lol ok Rainbow_Dashie

OpenStudy (anonymous):

i have already answer this one

OpenStudy (anonymous):

is that you Rainbow_Dashie?

OpenStudy (anonymous):

i was born the same year by the way Rainbow_Dashie

OpenStudy (anonymous):

but im in 8th

OpenStudy (anonymous):

yeah but i need another one